TEX and L^{A}TEX Macros for Linguistics

The following list is really a “work-in-progress”; it can’t claim to be exhaustive or complete, but is based on information which has come to light as people have posted message to the ling-tex list.

The purpose of ling-tex and of this partial listing of macros is to identify material which is available, and also stimulate testing, improvements to code and documentation, and so on, all with as much cooperation and assistance from the original authors as possible.

The material is all public domain, but the usual requests for citing authorship, not changing the contents without changing the file name, and so on also apply. These are the results of volunteers efforts, and a desire to share with others that work; this should always be kept in mind. Constructive criticism, helpful suggestions, offers of revised coding/wording, these are always more productive.

1 Index of .sty files listed

1. avm-doc.tex, avm.sty

2. cgloss4e.sty

3. chomsky.sty

4. cjl-glosses.tex

5. cm-lingmacros.sty

6. covingtn.tex covingtn.sty

7. glex.sty

8. gloss.tex, gloss.doc.

9. lingmacros.sty

10. numquote.doc, numquote.tex, enum.sty

11. pstrees

12. pstricks

13. tree-dvips

14. voorbeeldom.sty

2 Details on various .sty files

avm-doc.tex, avm.sty: By Christopher Manning (manning@csli.stanford.edu).

Macros for attribute-value matrices. Documentation available (but not printed in this collection).

CSLI: pub/TeXfiles.

cgloss4e.sty: Written by Hap Kolb and Craig Thiersch, this is a modified version of

covingtn.sty.

For glosses; as with covingtn.sty, doesn’t require ampersands (&) to align sets of glossed items.

“Following borrows from Covington’s style files inspired by Midnight by M. de Groot, adapted to be used with gbt3.sty: examples beginning with \ex can contain glosses directly. Default is Linguistic Inquiry style with all lines in \rm.”

No documentation, but file is heavily commented.

Posted to ling-tex; not currently available on archives.

chomsky.sty: Created by Michael Barr (barr@ ... ).

No documentation; however, file is heavily annotated. Some draft documentation has been written by Ch. Thiele

CTAN: tex-archive/macros/latex/contrib/misc/.

cjl-glosses.tex: Created by Michael Dunleavy, Carleton University; maintained by Ch.

Thiele (cthiele@ccs.carleton.ca).

Macros for glosses (seems to work in both plain TEX and in L^{A}TEX). Variants for centred, flush right or left glosses, and others. Some documentation; needs testing before it can be put out on the archives.

Posted to ling-tex list; not currently available on archives.

cm-lingmacros.sty: Modified version of Emma Pease’s lingmacros.sty, by Christopher Manning and Avery Andrews (Avery.Andrews@anu.edu.au).

CSLI: pub/TeXfiles.

covingtn.tex covingtn.sty: Created by Michael Covington.

L^{A}TEX macros for numbered examples, glosses, phrase structure rules, feature structures, discourse representation structures, exercises, reference lists, and miscellany. Documentation.

CTAN: tex-archive/macros/latex/contrib/covington/.

glex.sty: Created by Rob Norris; notes from Chet Creider.

L^{A}TEX macros for numbered glosses. All three lines of a gloss are input; by contrast, cjl-glosses.tex only takes care of the first 2 lines, requiring the 3rd line, the translation, to be formatted independently. On the other hand, glex.sty works with tabs, while cjl-glosses.tex groups each set of word-1 over gloss-1 within braces.

[Availability not yet determined.]

gloss.tex, gloss.doc: Part of the Midnight Macros set by Marcel van der Goot

(marcel@cs.caltech.edu).

Macros for vertically aligning words in consecutive sentences. Documentation.

CTAN: tex-archive/macros/macros/generic/midnight/.

lingmacros.sty: By Emma Pease, CSLI, Stanford.

Macros for numbered examples, trees, AVM structures.

CSLI: pub/TeXfiles.

numquote.doc, numquote.tex, enum.sty: Bob Mercer, U. of Western Ontario.

L^{A}TEX macros for automatic numbering of examples. Documentation.

Posted to ling-tex list; not currently available on archives.

pstrees: From Avery Andrews; requires Emma Pease’s tree-dvips package.

“This package consists of a preprocessor and some macro-definitions, by which linguistics-style trees can be specified as convenient indented lists, with spacing and line-drawing done automatically.”

CSLI: pub/TeXfiles (a tar file).

[Note: pstrees contains files which do not include the prefix “pstrees”; rather, the main files are “trees.**”, which may cause confusion if one is not careful – Ch.

pstricks: Created by Timothy Van Zandt (tvz@Princeton.EDU).

This is an extensive collection of PostScript macros that is compatible with most TeX macro packages, including Plain TEX, L^{A}TEX, AMS-TEX and AMS-LaTeX. Included are macros for color, graphics, rotation, trees and overlays. “PSTricks puts the icing (PostScript) on your cake (TEX)!” Documentation.

CTAN: tex-archive/graphics/pstricks/.

tree-dvips: Created by Emma Pease (emma@csli.stanford.edu).

CSLI PostScript drawing macros. These macros were originally created to draw the lines between nodes in the trees created by the tree macros in lingmacros.sty. They will only work with dvips version 541 or later (by Tomas Rokicki available on labrea.stanford.edu) but can be easily modified to be used with earlier versions of dvips and slightly less easily modified for other dvi to PostScript convertors. Documentation. [Formerly known as tree.tex.]

CSLI: pub/TeXfiles (a tar file).

treetex Created by Anne Brueggemann-Klein and Derick Wood.

Extensive tree-drawing macro set. Documentation available. See also: A. Brueggeman-Klein and Derick Wood (1989) “Drawing trees nicely with TEX,” Electronic Publishing 2.2.101–115.

TEX and L^{A}TEX Macros for Linguistics 4

voorbeeldom.sty: Created by Werenfried Spit (spit@vm.ci.uv.es, spit@ific.uv.es). L^{A}TEX document-style option which defines an enumerate-like environment for typesetting linguistic examples. No documentation, but .sty file has commented examples.

CTAN: tex-archive/macros/latex/contrib/misc/voorbeelden.sty.

3 What’s available where

Most TEX-ware is available from the CTAN archives, which can be accessed via ftp:

     Aston, UK  ftp.tex.ac.uk         /pub/archive/
     SHSU, Texas  ftp.shsu.edu          /tex-archive/
     Stuttgart, Germany ftp.uni-stuttgart.de  /pub/tex/

The CTAN holdings are too numerous to include here. In addition to this site, there has been a long-standing ftp site at Stanford:

csli.stanford.edu      /pub/TeXfiles/

The Index for the directory shows the following:

Phonetic.tar.Z: code for Phonetic Characters including some uppercase

Hausa. arm.tar.gz: Armenian 300pk and tfm fonts (old metafont) cslibib.bst.Z: CSLI bibliography style

1. citations in text look like (Barwise 1987) or (Barwise and Perry 1986) or (Barwise et al. 1985)

2. There are no labels in the bibliography (when used with acslibib.sty)

3. The first author’s name is reversed: last, first initial, first initial last. (adapted

        from anatsci.bst)

4. letters a, b, … are added to the year for multiple entries in label and reference list

5. A shortcite also exists (year only) (adapted from aaai-named.bst) cslibib.sty.Z: endnotes.sty: Endnotes instead of footnotes lingmacros.sty: Various linguistic macros. Well commented. permil.mf, permil.300gf, permit.tfm: code for a per mill sign pstrees.tar.Z: Avery Andrew’s tree macros suthesis.sty:] Stanford University thesis style tree.tar.Z: Macros for using PostScript to draw lines for trees (and other things).